OpenStudy (anonymous):

Can someone HELP me with this equation? -10<=3x-9<=-2

OpenStudy (anonymous):

x = 0, 1, 2 -1 <= 3x <= 7 -1/3 <= x <= 7/2

OpenStudy (anonymous):

3x-9 >= -10 x >= -1/3 3x-9 <= -2 x <= -7/3 -1/3 <= x <= 7/3

OpenStudy (anonymous):

pk51 made a human error: 7/3 not 7/2